# OBJEX(MT): Objective Extraction and Metacognitive Calibration for LLM-as-a-Judge under Multi-Turn Jailbreaks

### v1

@@ 5 Conclusion
Summary.
We introduced OBJEX(MT), a benchmark for testing whether an LLM can (i) recover the base objective of a multi-turn jailbreak and (ii) calibrate its confidence. A single human-aligned threshold calibrated on N=100N{=}100 items ($\tau^{\star}{=}0.61$) was frozen for all analyses. Under this regime, claude-sonnet-4 achieved the highest accuracy ($0.515$) and best metacognition (ECE $0.296$, Brier $0.324$, AURC $0.406$); gpt_4.1 and Qwen3-235B-A22B-FP8 were statistically tied in accuracy. Performance varied widely by source ($\sim\!0.167$$0.865$). Notably, less-accurate models reported high confidence (e.g., mean $\approx 0.88$ vs. accuracy $\approx 0.44$) and showed large Wrong@High-Conf rates near or above $50\%$ at $0.90$, indicating miscalibrated certainty.
Implications for LLM-as-a-Judge.
Judges should not be assumed to reliably infer latent objectives from long, noisy, adversarial dialogues. When feasible, provide the evaluation objective rather than asking the judge to infer it, and operationalize selective prediction/abstention to avoid trusting raw verbalized confidence. Reported evaluator quality can depend strongly on conversation structure; deployments should account for this sensitivity.
@@ 6 Limitations
Human supervision is limited to N=100N{=}100 items used to learn a single similarity threshold. This keeps cost low but introduces sampling variance and potential under-coverage of tactics and sources.
@@
We use gpt_4.1 as the sole similarity evaluator, which overlaps with one evaluated model. Although gpt_4.1 is not the top-performing extractor and model ordering is preserved under a categorical robustness check, evaluator–evaluatee affinity could still bias absolute scores.
@@
A single LLM judge may propagate idiosyncrasies (e.g., sensitivity to granularity or phrasing). We do not ensemble judges or combine LLM judging with complementary automated signals.
@@
We evaluate four sources without schema unification and construct the N=100N{=}100 calibration set with a fixed recipe that excludes harmless CoSafe items only for calibration. This can induce a calibration–evaluation distribution shift.
@@
For scoring, objective extraction is constrained to a single-sentence imperative base_prompt. This “flattens” multi-step or multi-objective attacks and may under-represent inherently compositional goals.
@@
Each extractor is run once with provider defaults (no replicas or self-consistency), isolating intrinsic behavior but potentially understating best-achievable accuracy under stronger prompting/decoding.
@@
We rely on verbalized confidence elicited by a fixed instruction. Self-reports are prompt-sensitive and not necessarily commensurate across models; we do not perform post-hoc calibration.
@@
Experiments are English, text-only; transfer to code-mixed, non-English, or multimodal jailbreaks is untested. Gold base_prompt strings come from source datasets; any label noise or granularity mismatch directly affects measured similarity.
@@ 7 Future Work
Expand beyond N=100N{=}100 with stratification by dataset, tactic, and length; include harmless items from all sources; report inter-annotator agreement and use aggregation models to capture uncertainty.
@@
Adopt heterogeneous multi-judge ensembles, hybrid scoring (e.g., bidirectional textual entailment, paraphrase checks, embedding similarity), and post-hoc calibration (e.g., isotonic). Use leave-one-family-out protocols to avoid evaluator overlap with tested models.
@@
Generalize the target to structured fields (action/target/constraints/success) or to a set of plausible objectives with confidences, enabling set-coverage metrics and finer error analysis.
@@
Evaluate selective prediction with explicit abstain/deferral policies, adaptive/binless ECE, and per-source risk–coverage curves; test stability of confidence under harmless paraphrases, re-orderings, and distractor insertions.
@@
Construct counterfactuals that vary dialogue length, evidence position, distractor density, and tactic composition to localize long-context and distractor-induced failures.
Broader settings and operations.
Extend to cross-lingual and multimodal jailbreaks; translate diagnostics into deployable policies (when to supply objectives to judges, when to escalate to humans, and how to set coverage targets) and provide lightweight recipes/scripts to reproduce calibration under alternative judges.
@@ B.2 Calibration details
  • Per-source counts: SafeMTData_1K 56, MHJ_local 22, SafeMTData_Attack600 22 (CoSafe absent).
@@
Sweeping $\{0.00,0.01,\dots,1.00\}$, F1 peaks at $\mathbf{\tau^{\star}{=}0.61}$ with F1=0.826, Precision=0.717, Recall=0.974. A bootstrap with B=1000B{=}1000 yields a 95% CI of [0.450, 0.900] for $\tau^{\star}$. [^1] Provenance: values from the released calibration.json and the Labeling sheet. The JSON’s N_total field reflects a metadata bug; the analysis uses N=100N{=}100.

### v2

@@ 5 Conclusion
ObjexMT tests whether an LLM can (i) recover the base objective of a multi-turn jailbreak and (ii) calibrate its confidence. Using a single human-aligned threshold calibrated on N=100N{=}100 items ($\tau^{\star}{=}0.61$) and then frozen, claude-sonnet-4 achieves the highest accuracy (0.515) and best calibration (ECE 0.296, Brier 0.324, AURC 0.406), while gpt-4.1 and Qwen3-235B-A22B-FP8 tie at 0.441 accuracy and display pronounced overconfidence (mean $\approx\!0.88$ vs. accuracy $\approx\!0.44$; Wrong@0.90 $\approx\!48$$52\%$). Performance varies widely across datasets ($\sim\!0.167$$0.865$), with MHJ comparatively easy and SafeMTData_Attack600/CoSafe challenging.
Implications for LLM-as-a-Judge. Judges should not be assumed to infer latent objectives from long, noisy, adversarial dialogues. When feasible, provide the evaluation objective and use selective prediction/abstention; evaluator quality is sensitive to conversation structure and should be accounted for in deployment.
Availability. The complete experimental dataset and accompanying templates are available at https://github.com/hyunjun1121/ObjexMT_dataset.
@@ 6 Limitations
Human-labeled calibration is limited to N=100N{=}100 to learn a single threshold, trading cost for sampling variance and possible under-coverage.
@@
A single evaluator (gpt-4.1) overlaps with one evaluated model; ordering is preserved under robustness checks, but affinity may bias absolute scores.
@@
A single LLM judge may propagate idiosyncrasies; we do not ensemble judges or add complementary automated signals.
@@
Dataset heterogeneity and the calibration recipe (excluding harmless CoSafe during calibration) can induce a calibration–evaluation shift.
@@
Constraining extraction to a single-sentence base_prompt may flatten multi-step or multi-objective attacks.
@@
Single-pass decoding with provider defaults isolates intrinsic behavior but may understate best-achievable accuracy.
@@
Verbalized confidence is prompt-sensitive and not necessarily commensurate across models; no post-hoc calibration is applied.
@@
Released artifacts include adversarial prompts and model outputs that may contain harmful strings. We filter for PII where feasible and preserve upstream dataset licenses; users should handle the data responsibly and adhere to research-only usage where applicable.
@@ 7 Future Work
Expand beyond N=100N{=}100 with stratified sampling (including harmless items), report inter-annotator agreement, and aggregate labels to capture uncertainty.
@@
Adopt heterogeneous multi-judge ensembles and hybrid scoring, apply post-hoc calibration, and use leave-one-family-out protocols to avoid evaluator overlap.
@@
Generalize objectives to structured fields or set-valued targets with confidences to enable set-coverage metrics and finer error analysis.
@@
Evaluate explicit abstain/deferral, adaptive/binless ECE, per-source risk–coverage, and confidence stability under benign paraphrases, re-orderings, and distractor insertions.
@@
Construct counterfactuals varying dialogue length, evidence position, distractor density, and tactic composition.
@@ B.2 Calibration details
  • Per-source counts: SafeMTData_1K 56, MHJ 22, SafeMTData_Attack600 22 (CoSafe absent).
@@
Sweeping $\{0.00,0.01,\dots,1.00\}$, $F_{1}$ peaks at $\mathbf{\tau^{\star}{=}0.61}$ with $F_{1}$=0.826, Precision=0.717, Recall=0.974. A bootstrap with B=1000B{=}1000 yields a 95% CI of [0.450, 0.900] for $\tau^{\star}$. [^2] Provenance: values from the released calibration.json and the Labeling sheet. The JSON’s N_total field reflects a metadata bug; the analysis uses N=100N{=}100.